OpenStudy (anonymous):

solve this equation x-y=12 x+y=9

OpenStudy (agreene):

\[(\frac{21}{2},-\frac{3}{2})\]

OpenStudy (anonymous):

so thats the answer thanks

OpenStudy (anonymous):

X=21/2 y=-3/2

OpenStudy (anonymous):

thanks DDgirl24

OpenStudy (anonymous):

Your welcome!:)

OpenStudy (anonymous):

thanks DDgirl24 and luis rivera